OpenStudy (anonymous):

WHAT ARE THE FACTORS OF THE FOLLOWING TRINOMIAL? 36m^2 + 24m +4

OpenStudy (anonymous):

\[36m^2+24m+4=4 (3 m+1)^2 \]

OpenStudy (anonymous):

the answer cant start like taht

OpenStudy (anonymous):

\[36 m^2+24 m+4=4 \left(9 m^2+6 m+1\right)=4 (3 m+1)^2 \]

OpenStudy (anonymous):

A (9m+24) (4m+2) B (9m +4) (4m+1) C (6M +2) ( 6M+2) D (6m + 4) (6m +1)

OpenStudy (anonymous):

C is the answer from your list presented. The following is your list expanded :\[\left\{36 m^2+114 m+48,36 m^2+25 m+4,36 M^2+24 M+4,36 m^2+30 m+4\right\} \]

OpenStudy (anonymous):

REALLY THATS THE ANSWER

OpenStudy (anonymous):

i mean polynomial

OpenStudy (anonymous):

(6M +2) ( 6M+2) =2(3M+1)*2(3M+1)=2*2*(3M+1)*(3M+1)=4*(3M+1)^2 The result above is the answer "completely factored".